# A Characterization for Tightness of the Sparse Moment-SOS Hierarchy

### v1

@@ 1. Introduction
(1.1) $\left\{\begin{array}[]{cl}\displaystyle\min_{x\in\mathbb{R}^{n}}&f(x)\coloneqq f_{1}(x_{\Delta_{1}})+\dots+f_{m}(x_{\Delta_{m}})\\ \mbox{s.t.}&h_{i}(x_{\Delta_{i}})=0,\,g_{i}(x_{\Delta_{i}})\geq 0,\,i=1,\ldots,m.\\ \end{array}\right.$
In the above, each $f_{i}$ is a polynomial and $h_{i},g_{i}$ are vectors of polynomials in $x_{\Delta_{i}}$. Throughout the paper, a minimizer for (1.1) means it is a global minimizer. We denote by $f_{\min}$ the minimum value of (1.1) and denote
@@
For a relaxation order $k$, we consider the sparse SOS relaxation
(1.3) $\left\{\begin{array}[]{cl}\max&\gamma\\ \mbox{s.t.}&f-\gamma\in\operatorname{Ideal}[h]_{spa,2k}+\operatorname{QM}[g]_{spa,2k}.\end{array}\right.$
The dual problem of (1.3) is the sparse moment relaxation
(1.4) $\left\{\begin{array}[]{cl}\min&\langle f,y\rangle\coloneqq\langle f_{1},y_{\Delta_{1}}\rangle+\cdots+\langle f_{m},y_{\Delta_{m}}\rangle\\ \mbox{s.t.}&\mathscr{V}_{h_{i}}^{spa,2k}[y]=0,\,L_{g_{i}}^{spa,k}[y]\succeq 0\ (i\in[m]),\\ &M_{spa}^{(k)}[y]\succeq 0,\,y_{0}=1,\,y\in\mathbb{R}^{\mathbb{U}_{k}}.\end{array}\right.$
We refer to Section 2 for the meaning of above notation. The sequence of relaxations (1.3)-(1.4) is called the sparse Moment-SOS hierarchy. The sparse relaxation has positive semidefinite (psd) matrix constraints whose sizes are much smaller than those of the dense relaxation. The sets $\Delta_{1},\ldots,\Delta_{m}$ are said to satisfy the running intersection property (RIP) if for every $j=2,\ldots,m$, it holds
@@
for some $t\leq j-1$. Let $f_{k}^{spa}$ and $f_{k}^{smo}$ denote the minimum values of (1.3) and (1.4) respectively. When RIP holds and each $\operatorname{Ideal}_{\Delta_{i}}[h_{i}]+\operatorname{QM}_{\Delta_{i}}[g_{i}]$ is archimedean, the sequence of $f_{k}^{spa}$ converges to the minimum value $f_{\min}$ of (1.1) asymptotically [15]. A convergence rate of the sparse Moment-SOS hierarchy is given in the recent work [13]. We refer to [8, 12, 15, 22, 31, 32, 33] for related work on the sparse polynomial optimization. Recently, sparsity has also been exploited to solve noncommutative polynomial optimization [11, 34]. Some applications can be found in [26, 38, 39, 40, 41]. Sparse Moment-SOS relaxations can be implemented in the software TSSOS [21, 35, 37].
In practice, people often observe that the sparse Moment-SOS hierarchy has finite convergence, i.e., it is tight. However, there exist examples for which the dense hierarchy is tight while the sparse one is not (see [31]). To the best of the authors’ knowledge, there is very few work to characterize when the sparse Moment-SOS hierarchy is tight.
@@
This paper characterizes tightness of the sparse Moment-SOS hierarchy (1.3)-(1.4). Our main contributions are:
  • We give a sufficient and necessary condition for the sparse Moment-SOS hierarchy to be tight. More precisely, when the optimal value of (1.3) is achievable, we show that $f_{k}^{spa}=f_{\min}$ if and only if there exist polynomials $p_{i}\in\mathbb{R}[x_{\Delta_{i}}]$ such that
@@
This paper is organized as follows. Some basics on polynomial optimization and algebraic geometry are reviewed in Section 2. In Section 3, we give a characterization for the tightness of sparse Moment-SOS hierarchy, and we study how to certify the tightness and get minimizers. Section 4 gives sufficient conditions for the sparse Moment-SOS hierarchy to be tight. The tightness of Schmüdgen type sparse Moment-SOS relaxations is investigated in Section 5. Some numerical experiments are presented in Section 6. Section 7 draws conclusions and makes some discussions.
@@ 2. Preliminaries
Notation
Denote by $\mathbb{N}$ the set of nonnegative integers and $\mathbb{R}$ the real field. For a positive integer $k$, let $[k]\coloneqq\{1,\ldots,k\}$. For a subset $\Delta_{i}\subseteq[n]$, denote by $\mathbb{R}^{\Delta_{i}}$ the space of real vectors $x_{\Delta_{i}}$. Let $\mathfrak{p}_{i}$ denote the projection from $\mathbb{R}^{n}$ to $\mathbb{R}^{\Delta_{i}}$ such that $\mathfrak{p}_{i}(x)=x_{\Delta_{i}}$ for all $x\in\mathbb{R}^{n}$. For $i,j\in[n]$, denote $\Delta_{ij}\coloneqq\Delta_{i}\cap\Delta_{j}$ and the projection $\mathfrak{p}_{ij}:\mathbb{R}^{\Delta_{i}}\rightarrow\mathbb{R}^{\Delta_{ij}}$ be such that $\mathfrak{p}_{ij}(x_{\Delta_{i}})=x_{\Delta_{ij}}$. The ring of polynomials in $x=(x_{1},\ldots,x_{n})$ with real coefficients is denoted as $\mathbb{R}[x]$. For a scalar or vector $p$ of polynomials, $\deg(p)$ denotes the maximal degree of its terms. For a power $\alpha=(\alpha_{1},\ldots,\alpha_{n})$, denote the monomial $x^{\alpha}\coloneqq x_{1}^{\alpha_{1}}\cdots x_{n}^{\alpha_{n}}$. For a degree $d$, the subset of polynomials in $\mathbb{R}[x]$ with degrees at most $d$ is denoted as $\mathbb{R}[x]_{d}$. The notation $\mathbb{R}[x_{\Delta_{i}}]$ and $\mathbb{R}[x_{\Delta_{i}}]_{d}$ are defined similarly. For $\phi\in\mathbb{R}[x]$ and subsets $A,B\subseteq\mathbb{R}[x]$, denote
@@
For $f\in\mathbb{R}[x]$, $\nabla f$ denotes its gradient with respect to $x$, and $\nabla_{x_{\Delta_{i}}}f$ denotes the gradient with respect to $x_{\Delta_{i}}$. The Hessian matrices $\nabla^{2}f$ and $\nabla^{2}_{x_{\Delta_{i}}}f$ are defined similarly. The Euclidean norm of $x$ is $\|x\|\coloneqq(|x_{1}|^{2}+\cdots+|x_{n}|^{2})^{1/2}$. Denote by $e_{i}$ the canonical basis vector such that the $i$th entry is $1$ and 00 otherwise. Let $\mathcal{S}^{n}$ be the space of all $n\times n$ real symmetric matrices. For $X\in\mathcal{S}^{n}$, $X\succeq 0$ (resp., $X\succ 0$) means $X$ is positive semidefinite (resp., positive definite).
@@ 2.1. Ideals and SOS polynomials
The vanishing ideal of a subset $V_{i}\subseteq\mathbb{R}^{\Delta{i}}$ is
@@
The notation $\Sigma[x]$ and $\Sigma[x]_{2k}$ are similarly defined. A tuple $g_{i}\coloneqq(g_{i,1},\ldots,g_{i,s_{i}})$ of polynomials in $\mathbb{R}[x_{\Delta_{i}}]$ generates the quadratic modules in $\mathbb{R}[x_{\Delta_{i}}]$ and $\mathbb{R}[x]$ respectively as
@@ 2.2. Sparse moments
Denote by $\mathbb{R}^{\mathbb{N}_{d}^{\Delta_{i}}}$ the space of all real vectors labeled by $\alpha\in\mathbb{N}_{d}^{\Delta_{i}}$. A vector $y_{\Delta_{i}}$ in $\mathbb{R}^{\mathbb{N}_{d}^{\Delta_{i}}}$ is called a truncated multi-sequence (tms) of degree $d$. For a given $y_{\Delta_{i}}\in\mathbb{R}^{\mathbb{N}_{d}^{\Delta_{i}}}$, the Riesz functional generated by $y_{\Delta_{i}}$ is the linear functional $\mathscr{R}_{y_{\Delta_{i}}}$ acting on $\mathbb{R}[x_{\Delta_{i}}]$ such that
$\mathscr{R}_{y_{\Delta_{i}}}(x^{\alpha})\,=\,(y_{\Delta_{i}})_{\alpha}\quad\mbox{for each}\quad\alpha\in\mathbb{N}_{d}^{\Delta_{i}}.$
@@
(2.9) $\langle p,y_{\Delta_{i}}\rangle\coloneqq\mathscr{R}_{y_{\Delta_{i}}}(p).$
@@
$\begin{array}[]{rcl}\mathscr{V}_{p}^{\Delta_{i},2k}[y_{\Delta_{i}}]&\coloneqq&\mathscr{R}_{y_{\Delta_{i}}}\left(p(x_{\Delta_{i}})[x_{\Delta_{i}}]_{k_{1}}\right),\\ L_{p}^{\Delta_{i},k}[y_{\Delta_{i}}]&\coloneqq&\mathscr{R}_{y_{\Delta_{i}}}\left(p(x_{\Delta_{i}})[x_{\Delta_{i}}]_{k_{2}}[x_{\Delta_{i}}]^{T}_{k_{2}}\right).\end{array}$
In the above, the Riesz functional is applied entry-wise and
@@
$\begin{array}[]{rcl}\mathscr{V}_{h_{i}}^{\Delta_{i},2k}[y_{\Delta_{i}}]&\coloneqq&(\mathscr{V}_{h_{i,1}}^{\Delta_{i},2k}[y_{\Delta_{i}}],\ldots,\mathscr{V}_{h_{i,\ell_{i}}}^{\Delta_{i},2k}[y_{\Delta_{i}}]),\\ L_{g_{i}}^{\Delta_{i},k}[y_{\Delta_{i}}]&\coloneqq&\mbox{diag}\left(L_{g_{i,1}}^{\Delta_{i},k}[y_{\Delta_{i}}],\ldots,L_{g_{i,s_{i}}}^{\Delta_{i},k}[y_{\Delta_{i}}]\right).\end{array}$
@@
(2.14) $\left\{\begin{array}[]{rcl}\mathscr{V}_{h_{i}}^{spa,2k}[y]&\coloneqq&\mathscr{V}_{h_{i}}^{\Delta_{i},2k}[y_{\Delta_{i}}],\\ L_{g_{i}}^{spa,k}[y]&\coloneqq&L_{g_{i}}^{\Delta_{i},k}[y_{\Delta_{i}}],\\ M^{(k)}_{spa}[y]&\coloneqq&\mbox{diag}\left(M_{\Delta_{1}}^{(k)}[y_{\Delta_{1}}],\ldots,M_{\Delta_{m}}^{(k)}[y_{\Delta_{m}}]\right).\end{array}\right.$
@@ 3. The characterization for tightness
(3.2) $\left\{\begin{array}[]{ccl}f_{k}^{spa}\coloneqq&\max&\gamma\\ &\mbox{s.t.}&f-\gamma\in\operatorname{Ideal}[h]_{spa,2k}+\operatorname{QM}[g]_{spa,2k}.\end{array}\right.$
@@
(3.3) $\left\{\begin{array}[]{rcl}f_{k}^{smo}\coloneqq&\min&\langle f,y\rangle=\langle f_{1},y_{\Delta_{1}}\rangle+\cdots+\langle f_{m},y_{\Delta_{m}}\rangle\\ &\mbox{s.t.}&\mathscr{V}_{h_{i}}^{spa,2k}[y]=0,L_{g_{i}}^{spa,k}[y]\succeq 0\ (i\in[m]),\\ &&M_{spa}^{(k)}[y]\succeq 0,y_{0}=1,\,y\in\mathbb{R}^{\mathbb{U}_{k}}.\end{array}\right.$

### v2

@@ 1. Introduction
(1.1) $\left\{\begin{array}[]{cl}\displaystyle\min_{x\in\mathbb{R}^{n}}&f(x)\coloneqq f_{1}(x_{\Delta_{1}})+\dots+f_{m}(x_{\Delta_{m}})\\ \mathrm{s.t.}&h_{i}(x_{\Delta_{i}})=0,\,g_{i}(x_{\Delta_{i}})\geq 0,\,i=1,\ldots,m.\\ \end{array}\right.$
In the above, each $f_{i}$ is a polynomial and $h_{i},g_{i}$ are vectors of polynomials in $x_{\Delta_{i}}$. We remark that for each $i$, the dimensions for the polynomial vectors $h_{i}$ and $g_{i}$ are not necessarily equal. It is also possible that $h_{i}$ or $g_{i}$ does not appear for some $i$. Throughout the paper, a minimizer for (1.1) means it is a global minimizer. We denote by $f_{\min}$ the minimum value of (1.1) and denote
@@
In this paper, we consider the sparse Moment-SOS hierarchy of semidefinite relaxations for solving sparse polynomial optimization. We refer to (3.2)-(3.3) in Section 3 for the exact formulations of sparse Moment-SOS relaxations. The sparse relaxation has positive semidefinite (psd) matrix constraints whose sizes are much smaller than those of the dense relaxation. The sets $\Delta_{1},\ldots,\Delta_{m}$ are said to satisfy the running intersection property (RIP) if for every $j=2,\ldots,m$, it holds
@@
for some $t\leq j-1$. The geometric meaning of RIP is as follows. Let $G=(V,E)$ be the graph associated with the correlative sparsity pattern $(\Delta_{1},\ldots,\Delta_{m})$, i.e., $V=\{1,\ldots,n\}$ and $(k_{1},k_{2})\in E$ if and only if $\{k_{1},k_{2}\}\subseteq\Delta_{i}$ for some $i$. The RIP is equivalent to that the sparsity pattern graph $G$ is chordal, i.e., all its cycles of length at least four have a chord (an edge is called a chord if it joins two non-adjacent nodes in the cycle). We refer to [2, 39] for more details on chordal graphs and the RIP. Let $f_{k}^{spa}$ and $f_{k}^{smo}$ denote the optimal values of (3.2) and (3.3) respectively. When the RIP holds and every $K_{\Delta_{i}}$ satisfies the archimedean condition, the sequence of $f_{k}^{spa}$ converges to the minimum value $f_{\min}$ of (1.1) asymptotically [13, 17]. A convergence rate of the sparse Moment-SOS hierarchy is given in the recent work [15]. We refer to [9, 14, 17, 24, 33, 34, 36] for related work on the sparse polynomial optimization. Recently, sparsity has also been exploited to solve noncommutative polynomial optimization [12, 37]. Some applications can be found in [28, 41, 42, 43, 44]. Sparse Moment-SOS relaxations can be implemented in the software TSSOS [23, 38, 40].
In practice, people often observe that the sparse Moment-SOS hierarchy has finite convergence, i.e., it is tight. However, there exist examples for which the dense hierarchy is tight while the sparse one is not (see [33]). To the best of the authors’ knowledge, there is very little work to characterize when the sparse Moment-SOS hierarchy is tight.
@@
This paper characterizes tightness of the sparse Moment-SOS hierarchy, which is given in (3.2)-(3.3). Our main contributions are:
  • We give a sufficient and necessary condition for the sparse Moment-SOS hierarchy to be tight. More precisely, when the optimal value of (3.2) is achievable, we show that $f_{k}^{spa}=f_{\min}$ if and only if there exist polynomials $p_{i}\in\mathbb{R}[x_{\Delta_{i}}]$ such that (see Section 2 for the meaning of notation below)
@@
  • We remark that this paper is the first work that characterizes tightness of the sparse Moment-SOS hierarchy. Also, we give a sparse version of the flat truncation condition to detect tightness and to extract minimizers for (1.1). It is generally hard to check tightness of the sparse Moment-SOS hierarchy [35]. There are almost no such results in the prior existing work, except for the case that moment matrices corresponding to intersections of blocks are rank one [17]. Moreover, we give several sufficient conditions for the sparse Moment-SOS hierarchy to be tight. For the dense case, the analogues of these sufficient conditions are studied in [5, 18, 30, 31]. However, the new sufficient conditions given in this paper are not studied in earlier work because the sparse Moment-SOS hierarchy requires additional assumptions for tightness. In Section 4, we show that these new sufficient conditions are satisfied for many cases; see, for instance, Theorems 4.2 and 4.6. We acknowledge that some of our results (e.g., Theorems 3.2, 3.5, 4.4, 4.5, 5.2) apply some techniques developed in the first author’s earlier work [29, 30, 31, 32].
This paper is organized as follows. Some basics on polynomial optimization and algebraic geometry are reviewed in Section 2. In Section 3, we give a characterization for the tightness of sparse Moment-SOS hierarchy, and we study how to certify the tightness and get minimizers. Section 4 gives sufficient conditions for the sparse Moment-SOS hierarchy to be tight. The tightness of Schmüdgen type sparse Moment-SOS relaxations is investigated in Section 5. Some numerical experiments are presented in Section 6. Proofs for some theorems in earlier sections are given in Section 7. Section 8 draws conclusions and makes some discussions.
@@ 2. Preliminaries
Notation Denote by $\mathbb{N}$ the set of nonnegative integers and $\mathbb{R}$ the real field. For a positive integer $k$, let $[k]\coloneqq\{1,\ldots,k\}$. For a subset $\Delta_{i}\subseteq[n]$, denote by $\mathbb{R}^{\Delta_{i}}$ the space of real vectors $x_{\Delta_{i}}$. Let $\mathfrak{p}_{i}$ denote the projection from $\mathbb{R}^{n}$ to $\mathbb{R}^{\Delta_{i}}$ such that $\mathfrak{p}_{i}(x)=x_{\Delta_{i}}$ for all $x\in\mathbb{R}^{n}$. For $i,j\in[n]$, denote $\Delta_{ij}\coloneqq\Delta_{i}\cap\Delta_{j}$ and the projection $\mathfrak{p}_{ij}:\mathbb{R}^{\Delta_{i}}\rightarrow\mathbb{R}^{\Delta_{ij}}$ be such that $\mathfrak{p}_{ij}(x_{\Delta_{i}})=x_{\Delta_{ij}}$. The ring of polynomials in $x=(x_{1},\ldots,x_{n})$ with real coefficients is denoted as $\mathbb{R}[x]$. For a scalar or vector $p$ of polynomials, $\deg(p)$ denotes the maximal degree of its terms. For a power $\alpha=(\alpha_{1},\ldots,\alpha_{n})$, denote the monomial $x^{\alpha}\coloneqq x_{1}^{\alpha_{1}}\cdots x_{n}^{\alpha_{n}}$. For a degree $d$, the subset of polynomials in $\mathbb{R}[x]$ with degrees at most $d$ is denoted as $\mathbb{R}[x]_{d}$. We define $\mathbb{R}[x_{\Delta_{i}}]$ and $\mathbb{R}[x_{\Delta_{i}}]_{d}$ similarly. For $\phi\in\mathbb{R}[x]$ and subsets $A,B\subseteq\mathbb{R}[x]$, denote
@@
For $f\in\mathbb{R}[x]$, $\nabla f$ denotes its gradient with respect to $x$, and $\nabla_{\Delta_{i}}f$ denotes the gradient with respect to $x_{\Delta_{i}}$. The Hessian matrices $\nabla^{2}f$ and $\nabla^{2}_{\Delta_{i}}f$ are defined similarly. The Euclidean norm of $x$ is $\|x\|\coloneqq(|x_{1}|^{2}+\cdots+|x_{n}|^{2})^{1/2}$. Denote by $e_{i}$ the canonical basis vector such that the $i$th entry is $1$ and 00 otherwise. Let $\mathcal{S}^{n}$ be the space of all $n\times n$ real symmetric matrices. For $X\in\mathcal{S}^{n}$, $X\succeq 0$ (resp., $X\succ 0$) means $X$ is positive semidefinite (resp., positive definite).
@@ 2.1. Ideals and SOS polynomials
The vanishing ideal of a subset $V_{i}\subseteq\mathbb{R}^{\Delta_{i}}$ is
@@
We define $\Sigma[x]$ and $\Sigma[x]_{2k}$ similarly. A tuple $g_{i}\coloneqq(g_{i,1},\ldots,g_{i,s_{i}})$ of polynomials in $\mathbb{R}[x_{\Delta_{i}}]$ generates the quadratic modules in $\mathbb{R}[x_{\Delta_{i}}]$ and $\mathbb{R}[x]$ respectively as
@@ 2.2. Sparse moments
Denote by $\mathbb{R}^{\mathbb{N}_{d}^{\Delta_{i}}}$ the space of all real vectors labeled by $\alpha\in\mathbb{N}_{d}^{\Delta_{i}}$. A vector $y_{\Delta_{i}}$ in $\mathbb{R}^{\mathbb{N}_{d}^{\Delta_{i}}}$ is called a truncated multi-sequence (tms) of degree $d$. For a given $y_{\Delta_{i}}\in\mathbb{R}^{\mathbb{N}_{d}^{\Delta_{i}}}$, the Riesz functional generated by $y_{\Delta_{i}}$ is the linear functional ${R}_{y_{\Delta_{i}}}$ acting on $\mathbb{R}[x_{\Delta_{i}}]$ such that
${R}_{y_{\Delta_{i}}}(x^{\alpha})\,=\,(y_{\Delta_{i}})_{\alpha}\quad\mbox{for each}\quad\alpha\in\mathbb{N}_{d}^{\Delta_{i}}.$
@@
(2.9) $\langle p,y_{\Delta_{i}}\rangle\coloneqq{R}_{y_{\Delta_{i}}}(p).$
@@
$\begin{array}[]{rcl}{V}_{p}^{\Delta_{i},2k}[y_{\Delta_{i}}]&\coloneqq&{R}_{y_{\Delta_{i}}}\left(p(x_{\Delta_{i}})[x_{\Delta_{i}}]_{k_{1}}\right),\\ L_{p}^{\Delta_{i},k}[y_{\Delta_{i}}]&\coloneqq&{R}_{y_{\Delta_{i}}}\left(p(x_{\Delta_{i}})[x_{\Delta_{i}}]_{k_{2}}[x_{\Delta_{i}}]^{T}_{k_{2}}\right).\end{array}$
In some literature, the localizing matrix $L_{p}^{\Delta_{i},k}[y_{\Delta_{i}}]$ is also denoted as $M_{k-\lceil\deg(p)/2\rceil}(py_{\Delta_{i}})$; see [17, 19, 24, 40]. In the above, the Riesz functional is applied entry-wise and
@@
$\begin{array}[]{rcl}{V}_{h_{i}}^{\Delta_{i},2k}[y_{\Delta_{i}}]&\coloneqq&({V}_{h_{i,1}}^{\Delta_{i},2k}[y_{\Delta_{i}}],\ldots,{V}_{h_{i,\ell_{i}}}^{\Delta_{i},2k}[y_{\Delta_{i}}]),\\ L_{g_{i}}^{\Delta_{i},k}[y_{\Delta_{i}}]&\coloneqq&\mbox{diag}\left(L_{g_{i,1}}^{\Delta_{i},k}[y_{\Delta_{i}}],\ldots,L_{g_{i,s_{i}}}^{\Delta_{i},k}[y_{\Delta_{i}}]\right).\end{array}$
@@
(2.14) $\left\{\begin{array}[]{rcl}{V}_{h_{i}}^{spa,2k}[y]&\coloneqq&{V}_{h_{i}}^{\Delta_{i},2k}[y_{\Delta_{i}}],\\ L_{g_{i}}^{spa,k}[y]&\coloneqq&L_{g_{i}}^{\Delta_{i},k}[y_{\Delta_{i}}],\\ M^{(k)}_{spa}[y]&\coloneqq&\mbox{diag}\left(M_{\Delta_{1}}^{(k)}[y_{\Delta_{1}}],\ldots,M_{\Delta_{m}}^{(k)}[y_{\Delta_{m}}]\right).\end{array}\right.$
@@ 3. The characterization for tightness
(3.2) $\left\{\begin{array}[]{ccl}f_{k}^{spa}\coloneqq&\max&\gamma\\ &\mathrm{s.t.}&f-\gamma\in\operatorname{Ideal}[h]_{spa,2k}+\operatorname{QM}[g]_{spa,2k}.\end{array}\right.$
@@
(3.3) $\left\{\begin{array}[]{rcl}f_{k}^{smo}\coloneqq&\min&\langle f,y\rangle=\langle f_{1},y_{\Delta_{1}}\rangle+\cdots+\langle f_{m},y_{\Delta_{m}}\rangle\\ &\mathrm{s.t.}&{V}_{h_{i}}^{spa,2k}[y]=0,L_{g_{i}}^{spa,k}[y]\succeq 0\ (i\in[m]),\\ &&M_{spa}^{(k)}[y]\succeq 0,y_{0}=1,\,y\in\mathbb{R}^{\mathbb{U}_{k}}.\end{array}\right.$